At t = 0. a rectangular coil of resistance R = 2 o
Answer #1

Here ,

W = 3 cm

L = 8 cm

B = 1.6 T

R = 2 Ohm

let the current in the loop is I

I = B * v * L/R

I = 1.6 * 0.05 * 0.03/2

I = 1.2 *10^-3 A

force required to maintain this velocity = B*I*L

force required to maintain this velocity = 1.2 *10^-3 * 1.6 * 0.02

force required to maintain this velocity = 3.84 *10^-5 N

the force required to maintain this velocity is 3.84 *10^-5 N